FEL Upgrade Project Weekly Brief –May 27-31, 2002




More than 80% of the support stands for electron beam transport hardware have been installed in the FEL vault. Three of the "girder assemblies" which contain recently released quadrupole magnets have been fully assembled and released for installation in the vault.


FEL project management and JLab management hosted a meeting of our "Maritime Technical Advisory Committee (MTAC) on May 28-29. MTAC has advised the lab on FEL development for Navy interests since the FEL program was formally kicked-off in 1996. MTAC is chaired by Admiral Al Baciocco, and includes Admiral Ted Parker, Admiral Ken Malley and the former director of the Naval Research Labs, Alan Berman, as members. We have always appreciated the wise and experienced council offered by MTAC members and were again grateful for their cogent and frank input. We expect a formal report from MTAC in the next two weeks.

On Wednesday May 29th, George Neil attended a meeting of the Navy R&D teams who are participating in the laser materials damage (lethality) studies funded by the DOD Joint Technology Office. Planning continued for the next phase of lethality measurements to be done using the 10 kW Upgrade along with our partners from NRL and MIT-Lincoln Labs.

The CRADA between JLab and AES, which was approved by DOE last week, was signed this week by Jefferson Lab and forwarded to AES.

The loan agreement between Cornell and Jefferson Lab for the loan of Cornell’s undulator to Jefferson Lab for use on the UV FEL Upgrade was signed and forwarded to Cornell.

WBS 4 (Injector):

The gun chamber was cleaned and mounted back on the stand, it is ready for installation in the vault. Design work continued on the SF6 can modifications. The ball cathode and shield door polishing are complete. Arcing occurred on the support tube while implanting which was most likely dust related. A HEPA filter unit is being installed to improve the particulate cleanliness around the implant chamber and gas line filters and CO2 dusters are being ordered. The shield door will be implanted next with the ball cathode to follow.



Gun HVPS - The delivery of the SF6 tank was moved back out to 6/16/02 by the vendor. They had a problem with some flanges they purchased for the tank. The first conditioning resistor was sprayed with the resistive ink and air dried. The incremental weight gain and film thickness checked out. It will be fired early next week.

WBS 6 (RF):

Quarter Klystrons - The 2nd and 3rd klystron carts are finished at the Machine Shop. They will be delivered and the klystrons installed early next week.

Quarter HVPS - New dual 18 volt power supplies were ordered for the klystron heater monitor circuits. The circuits were tested and found to be linear over the range of interest. The low level wiring is finished for one unit and nearly finished for the other. A 50 kW klystron is being connected as a test load for the HVPS. The PLC programming and any new wiring is the next step before testing can begin.

WBS 8 (Instrumentation):

The quadrupole girder assembly has moved into full swing. They are being assembled and delivered at 1 or 2 a day, depending on complexity. There are a few parts that are still trickling, like shielded pump drops, but this should not hold up assembly. Improvements continue to be made on the magnet test stand; the PCI based motor driver code is being done by Mike Necaise. This will begin the transition out of CAMAC to an entirely PC/LabView based test stand. This work had been scheduled for some time by Tom Hiatt and crew, but with the quads now moving on the decision was made to begin the conversion. Many thanks to the magnet test guys and the vacuum assembly crews!

Fabrication nearly complete for 3 of the 4 new OMS chassis. One OMS chassis has been installed in the control room for further electrical testing and the software installed for compatibility. The mechanical drawing for the BLM photo multiplier assembly has been revised and awaiting sign off. Design upgrade for the raster power supplies has begun. Maintenance for the LeCroy HV mainframe has been completed,tested and re-installed in the rack. Continuing cable clean up in the quarter cryo region. Website maintenance and upgrades are continuing. The schematic for the beam viewer Canbus interface has been finished and PCB layout has begun. The two additional student positions are on hold due to a hiring freeze. The timing, end of semester, could not be worse for trying to attract the brightest and most talented. We had been stringing along a prospect but he lost hope in JLab. Our coordinator Deborah, who has been out for two months, has had surgery and came through with flying colors. Her recovery prospects are very good, but her return to work will not be for a number of months. My planning and scheduling for the installation tasks are based on the additional students. By not having them I will be forced to seek higher cost help elsewhere, this will adversely affect my cost and schedule performance.









WBS 9 (Transport):


Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. We are studying the epoxy problems.

• Upper and lower core’s plates remain in final machining. We will check completed cores on June 7.

Injector Dipoles (DU/DV)

• At Master Machine, core pieces remain in final machining. All hardware and tooling is completed, first article will be assembled next week.

• At Wang NMR, they remain working on potting the coils.

Arc Dipoles (GY, GX, GQ)

• Process Equipment Co. (PECo) continued manufacturing the cores and parts of the GQ-GX dipoles. GQ cores are at heat treat. GX cores are in rough machining. Completed heat treating of GQ core material, started final machining. GX material is in heat treating process.

• Wang NMR continues working on the winding fixture for the GQ-GX coils.

• Bosma Machine is shipping the first article GY Core to WANG for assembly. We will inspect the second core’s gap on June 5th. Received RFI 007 with costs for approval. Shipped first article GY to Wang. Second core is being assembled. Spare steel will ship back to JLAB this week.


3 inch quad (QX)

• We found that the electric current set point reliability of the Trim Card Power Supplies is about ±3 parts in 1000, well below David Douglas’ specification for magnetic field set ability of 1 part per 1000. Grist for getting better trim supplies. After some down time on the magnet test stand, the full production retesting of the QX quads started.

• Production measurements have resumed and 6 QX quads have been released for installation. Improvements to stand precision are still being pursued.

• Final shipment of QX assemblies will be shipped next week.

Trim Quad (QT)

• Milhous Control of Virginia continues manufacturing the remaining quads.

• No word on the GC corrector because Tom Hiatt is consumed with test stand problems.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet. They are finishing up details. Details of SF are complete except for stand mounting. Coordinating with JLAB designer to complete. Preliminary stand/support design is done.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The procurement documents for the ARC Chambers are in composition.

• A number of stands were delivered and bolted into position.

• Work continues on the recirculation region in front of the Cryomodule.

• Design of the optical chicane region with its chamber continues.

• Design of the regions around the wiggler continues.